# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 1.1.2 Reweighted Spectral Partitioning
For a graph $G$ with maximum degree $\Delta$,
The “hard” direction of Theorem 1.6 essentially uses a (lossy) dimension-reduction step, which is also quite simple algorithmically. We note for later that there is also an easier dimension-reduction step that is useful in some cases for smaller initial dimension $d$: Specifically, it can be shown that $\gamma^{(1)}(G)\lesssim d\gamma^{(d)}(G)$ [OTZ22, Proof of Proposition 2.9].
Unlike $\lambda_{2}$, no non-trivial bounds on $\lambda_{2}^{*}$ are yet known, and hence reweighted spectral partitioning is not yet known to produce good separators. This is the primary focus of our work.
1.1.3 Other Methods
The guarantees obtained by spectral partitioning are of an unusual form compared to typical approximation algorithms. Some more typical results are known. In particular, for a general graph with $n$ vertices, and $m\geq n$ edges, an $\operatorname{\mathcal{O}}(m^{1+\varepsilon})$-time algorithm (for any $\varepsilon>0$) is known that can obtain an $\operatorname{\mathcal{O}}(\sqrt{\log n})$-approximation of either edge or vertex expansion [LTW24]. However, it is quite impractical due to the use of almost-linear time min-cost flow [CKL+22b, CKL+22a]. This result strengthened many other prior results obtaining slightly weaker bounds or slower algorithms, although a slightly more refined approximations is still possible in polynomial time for many cases [FHL05, FHL08] (we will briefly discuss this towards the end of our paper, since we will use some related techniques).
1.2 Results and Organization
We will give bounds on $\lambda_{2}^{*}$, $\gamma^{(1)}$, and $\gamma^{(d)}$ for a large variety of graph classes, summarized in Table 3 and Table 4. In turn, these imply bounds on (vertex) separator sizes, as well as bounds on the resulting separators from reweighted spectral partitioning, summarized in Table 5. This also slightly improves the dependence on dimension for separator bounds in some geometric graph classes. Most of our bounds on $\lambda_{2}^{*}$ will primarily adapt existing machinery of bounds for $\lambda_{2}$. However, for $g$ graphs, we are able to make some notable improvements to the techniques, and our bound on $\lambda_{2}^{*}$ also implies a new bound on $\lambda_{2}$. These techniques are the topic of Section 3.2. The main new tool is the use of something we call “uniform” shallow minors, although these are used in a very different manner compared to previous work leveraging forbidden shallow minors.
{TAB}
(2,1pt,1pt)[4pt]—c—c——c—c—c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(1)}\lesssim$
Planar $\dfrac{1}{n}$
[…]
$K_{h}$-minor-free $\dfrac{(h\log h\log\log h)^{2}}{n}$
RIG over $K_{h}$-minor-free $\dfrac{\Delta\,h^{6}(\log\log h)^{2}}{n}$
Table 3: Our asymptotic upper bounds on $\gamma^{(1)}$ and $\lambda_{2}^{*}$ for several graph classes given in Corollary 4.9, Theorem 3.5 and Theorem 3.17.
{TAB}
(2,1pt,1pt)[4pt]—c—c——c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(d)}\lesssim$
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\left(\frac{k}{n}\right)^{\frac{2}{d}}$
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\left(\frac{k}{n}\right)^{\frac{2}{d}}$
Table 4: Our asymptotic upper bounds on $\gamma^{(d)}$ and $\lambda_{2}^{*}$ for geometric graph classes, given in Section 3.1.
{TAB}
(3,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c—c—c— Graph class Separator size $\lesssim$ Poly-time separator size $\lesssim$
Planar $\sqrt{n}$ $\sqrt{n}$
[…]
$K_{h}$-minor-free $(h\log h\log\log h)\sqrt{n}$ $\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n}$
RIG over $K_{h}$-minor-free $h^{3}\log\log h\sqrt{n\Delta}$ $h^{3}\log\log h\sqrt{n\Delta\log\Delta}$
$k$-ply neighbour -hood system in $\mathbb{R}^{d}$ $\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d}$ $\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d}$
$k$-nearest-neigh -bour graph in $\mathbb{R}^{d}$ $\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d}$ $\sqrt{d+\log k}\cdot k^{1/d}n^{1-1/d}$
Table 5: From the bounds on $\lambda_{2}^{*}$ stated in Table 3 and Table 4, we obtain (asymptotic) separator sizes via Theorem 1.5, and polynomial-time computable (asymptotic) separator sizes obtainable via reweighted spectral partitioning as stated in Theorem 5.3.
Overall, the techniques for these bounds can be categorized into “geometric intersection graph” techniques, which we consider in Section 3, and “metric” techniques, which we consider in Section 4. These categories also apply to the existing bounds on $\lambda_{2}$, so we discuss the history of these bounds and techniques in Section 2.
In Section 5, we will give a refined Cheeger-style inequality for vertex expansion that can be specialized for a number of graph classes. This refinement will involve the “padded partition modulus”, which we will define in the next section. The most interesting specialized (constructive) implication of the yet-unstated refined Cheeger inequality is:
Corollary 1.7 (Refined Cheeger Inequality for Vertex Expansion in Minor-free Graphs).
For a graph $G$ with no $K_{h}$-minor,
$\frac{\psi(G)^{2}}{\min\{\log\Delta,(\log h)^{2}\}}\lesssim\lambda_{2}^{*}(G)\lesssim\psi(G).$
For the class of $K_{h}$-minor-free graphs, where $h$ is fixed (which includes a superclass of planar graphs), the dependence on the maximum degree $\Delta$ is completely eliminated. Hence, when combined with the upper bound on $\lambda_{2}^{*}$ for $K_{h}$-minor-free graphs, it is possible to use reweighted spectral partitioning (in polynomial time) to obtain separators with only a slightly sub-optimal dependence on $h$. Analogous forms of this statement for other graph classes are used to obtain the “poly-time separator size” column in Table 5.
In Section 6, we will give a high-level view displaying some surprising symmetries between quantities related to $\lambda_{2}^{*}$, and quantities used to obtain the “metric” bounds in Section 4, which suggest promising avenues for future work.
1.3 Parallel Work
In parallel to this work, Kam Chuen Tung independently derived the same bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$ for planar graphs stated in Theorem 3.5, as well as the bound on $\lambda_{2}^{*}$ for bounded-genus graphs stated in Corollary 4.9 (which forms one part of the bound given in Table 3), and a slightly weaker bound on $\lambda_{2}^{*}$ for $K_{h}$-minor-free graphs compared to the one stated in Corollary 4.9. The gap for forbidden-minor graphs is due to improvements external to both works. The technique for planar graphs is identical, while the techniques for the other two bounds differ very slightly (specifically, we bound $\gamma^{(1)}$ first, rather than bounding $\lambda_{2}^{*}$ directly). These results have appeared in his recent PhD thesis [Tun25, Chapter 7], which also includes some interesting generalizations. Both projects originated from simultaneous independent course projects in 2022.
2 Detailed Background
In this section, we give some more detailed background that is important to our methods. We will first discuss the history of bounds on $\lambda_{2}(G)$ for a few graph classes in more detail. Then we will discuss two groups of tools used for some of those bounds.
2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
Bounds on $\lambda_{2}(G)$ are known for a wide variety of graph classes. One of the first, and essentially the only to date that is known to be tight (up to constant factors), is the one for planar graphs, discovered by Spielman and Teng [ST96, ST07]:
Theorem 2.1 ([ST96, ST07, Theorem 3.3]).
For a planar graph $G$ of $n$ vertices and maximum degree $\Delta$, $\lambda_{2}(G)\leq\frac{8\Delta}{n}$.
Combined with Theorem 1.1, this obtains the weaker form of the planar separator theorem. Spielman and Teng also conjectured several bounds for $\lambda_{2}$ in several other classes of graphs, including bounded-genus graphs and minor-free graphs. Of particular note (since we will propose a generalized form) is their conjecture for bounded-genus graphs:
Conjecture 2.2 ([ST96, ST07, Conjecture 1]).
For a graph $G$ of $n$ vertices, genus $g$ and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta g}{n}$.
We use the notation $a\lesssim b$ throughout the paper to denote that $a\leq\operatorname{\mathcal{O}}(b)$.
Progress towards this conjecture has been the subject of several follow-up papers [Kel04, Kel06b, BLR10, KLPT09]. The conjecture was first partly answered by Kelner [Kel04, Kel06b], who showed the following:
Theorem 2.3 ([Kel04, Kel06b, Theorem 2.3], expanded proof in [Kel06a, Chapters 11 and 12]).
For a triangulated $g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
This result is of particular interest since the proof is entirely geometric, and it was essentially the last result of this kind to have a geometric proof. As mentioned earlier, the references for this result claim that it applies to general $g$ graphs. This is due to the fact that any $g$ graph can be triangulated by adding edges (i.e., without adding new vertices), and adding edges only increases $\lambda_{2}$. The issue with this logic is that adding edges naively could also very significantly increase $\Delta$, even to the order of $n$. A method of triangulation by adding edges that asymptotically preserves $\Delta$ is known for planar graphs [KB92, Kan93, KB97], but, to the best of our knowledge, no such result is known for $g$ graphs. We will later bypass this issue with a careful triangulation that also adds vertices.
Orthogonal progress on the conjecture was separately made by Biswal, Lee and Rao, who gave a bound polynomial in $g$ instead of $\Delta$, by making use of metric embeddings [BLR10]:
Theorem 2.4 ([BLR10, Theorem 5.2]).
For a graph $G$ of $n$ vertices with genus $g$, maximum degree $\Delta$, and at least $n\gtrsim g$ vertices, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta\cdot g^{3}}{n}$.
Note that the restriction of $n\gtrsim g$ is actually unnecessary: It follows from Geršgorin circle theorem [Ger31] that $\lambda_{2}(G)\leq 2\Delta$ for all graphs. So, if $n\lesssim g$, then $\lambda_{2}(G)\lesssim\Delta\lesssim\frac{\Delta g}{n}\lesssim\frac{\Delta\cdot g^{3}}{n}$.
Their methods also allowed them to give a bound for $K_{h}$-minor-free graphs:
Theorem 2.5 ([BLR10, Theorem 5.3]).
For a graph $G$ of $n$ vertices which contains no $K_{h}$ minor, has maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta h^{7}}{n}$. If $G$ has at least $n\gtrsim h\sqrt{\log h}$ vertices, then $\lambda_{2}(G)\lesssim\frac{\Delta\cdot h^{6}\log h}{n}$.
Note that the second of these bounds is stronger (and more generally applies without the restriction $n\gtrsim h\sqrt{\log h}$) by the same argument as above.
The bounds of Biswal, Lee, and Rao have since been slightly tightened due to improvements of some mathematical tools used to devise them, resulting in the improved bounds given in Table 2. In particular, one of the key steps in their method is to apply a certain form of metric embedding. This metric embedding induces a $g^{2}$ factor in the bounded-genus bound, and an (asymptotic) $h^{4}$ factor in both forbidden minor bounds. These metric embedding results have since been improved to $(\log g)^{2}$ [LS10] and $(\log h)^{2}$ [CF25], respectively. Note that the latter bound is more general. Moreover, the existing $\sqrt{\log h}$ factor in the vertex count requirement in the second forbidden minor bound (which appears squared in the corresponding bound on $\lambda_{2}$) arises from a weak form of Hadwiger’s conjecture, and has since been improved to $\log\log h$ [DP25]. We discuss these improvements in a bit more detail later in this section.
A similar theorem is also known for region intersection graphs:
Theorem 2.6 ([Lee17]).
Let $G$ be a $K_{h}$-minor-free graph. For a graph $\hat{G}\in\text{rig}(G)$ with maximum degree $\Delta$ and $n$ vertices,
$\lambda_{2}(G)\lesssim\frac{\Delta^{2}\cdot h^{6}\log h}{n}.$
Compared to the bounds for minor-free and bounded-genus graphs, this bound for region intersection graphs is weaker, and does not generalize the bound for minor-free graphs (even for constant $h$) as we might expect. To do so for constant $h$ would require replacing one of the $\Delta$ (the maximum degree) factors by $m/n$ (the average degree). Unfortunately, it is not clear if this could be done.
2.2 Extremal Spread
To obtain their bounds on $\lambda_{2}$, Biswal, Lee, and Rao [BLR10] studied the “$L_{2}$-extremal spread”. In slightly more generality:
@@ 2.2 Extremal Spread
$\overline{s}_{p}(G):=\sup_{\omega:V(G)\to\mathbb{R}_{\geq 0},||\omega||_{p}\leq 1}\sum_{u,v\in V(G)}d_{G}^{\omega}(u,v).$
{TAB}
(2,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c— Graph class $\overline{s}_{2}(G)\gtrsim$ Refs.
@@
$K_{h}$-minor-free $n\geq 4h$ $\dfrac{n^{2}}{h^{3/2}}$ [BLR10, Theorems 3.9 and 3.11]
$K_{h}$-minor-free $n\gtrsim h\sqrt{\log h}$ $\dfrac{n^{2}}{h\sqrt{\log h}}$ [BLR10, Theorems 3.9 and 3.11]
RIG over $K_{h}$-minor-free $\dfrac{n^{2}}{h\sqrt{\Delta\log h}}$ [Lee17, Corollary 5.4]
Table 6: Asymptotic lower bounds on the $L^{2}$-extremal spread $\overline{s}_{2}(G)$ for various graph classes. $n$ is always the number of vertices, and $\Delta$ is the maximum degree.
Note that, by definition, $\overline{s}_{2}(G)\geq\overline{s}_{1}(G)$. A number of bounds for $\overline{s}_{2}$ are known, which we will review. A summary can be found in Table 6.
By studying the dual of the $L_{2}$-extremal spread, and applying crossing number bounds, Biswal, Lee, and Rao were able to prove the following results for special graph classes:
Theorem 2.8 ([BLR10, Theorem 3.1]).
For a $g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
@@
Theorem 2.9 ([BLR10, Theorems 3.9 and 3.11]).
For a $K_{h}$-minor-free graph $G$ on $n$ vertices, where $n\geq 4h$,
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h^{\frac{3}{2}}}.$
Furthermore, if $n\gtrsim h\sqrt{\log h}$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h\sqrt{\log h}}.$
The bound for bounded-genus graphs is essentially optimal, while the second bound for $K_{h}$-minor-free graphs is essentially optimal up to the $\sqrt{\log h}$ factor. Two recent results actually improve the $\sqrt{\log h}$ in the second bound: Norin, Postle, and Song [NPS23] first showed that it may be replaced with $(\log h)^{\frac{1}{4}+\varepsilon}$, for any $\varepsilon>0$. Delcourt and Postle [DP25] further improved this to $\log\log h$. The possibility of reducing this factor to $\operatorname{\mathcal{O}}(1)$ is closely related to Hadwiger’s conjecture, which is a major open problem in graph theory [Die25].
For region intersection graphs, Lee [Lee17] instead studied the dual of the $L^{1}$-extremal spread, generalizing results of Matoušek for string graphs [Mat14]. In particular, Lee obtained the following result based on an argument of Matoušek’s:
Theorem 2.10 ([Lee17, Theorem 2.2]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$: If $G$ has no $K_{h}$-minor, then $\overline{s}_{1}(\hat{G})\gtrsim\frac{n^{2}}{h\sqrt{m\log h}}$.
Note that our definition of $L^{1}$-extremal spread differs from Lee’s by a factor of $n$. Using essentially the same argument, Lee also obtained a relatively weak bound on $\overline{s}_{2}$:
Theorem 2.11 ([Lee17, Corollary 5.4]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$, so that $\Delta$ is the maximum degree of $\hat{G}$: If $G$ has no $K_{h}$-minor, then $\overline{s}_{2}(\hat{G})\gtrsim\frac{n^{2}}{h\sqrt{\Delta\log h}}$.
Similarly to Theorem 2.9, the $\sqrt{\log h}$ factor in these two theorems can also be replaced by a $\log\log h$ factor.
2.3 Non-Expansive Metric Embeddings
In proving their respective bounds on $\lambda_{2}$, Biswal, Lee, and Rao [BLR10], and Lee and Sidiropoulos [LS10] both made use of a certain forms of metric embeddings as a key step. We discuss such embeddings here, along with some other relevant literature. Importantly, this is the step that produces most of the loss for the bounds on $\lambda_{2}$.

### v4

@@ 2.3 Bounds via Geometry with Ball-Intersection Graphs
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then,
$\gamma^{(1)}(G)\lesssim\gamma^{(3)}(G)\leq\frac{8}{n}.$
Proposition 2.5.
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Corollary 2.6.
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Proposition 2.7.
$\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$
Note that Proposition 2.4 implies the planar separator theorem when applied to Theorem 1.4, while Propositions 2.5 and 2.6 both imply the new separator theorems for their respective classes mentioned in Corollary 1.6.
Many of the techniques for obtaining these bounds will be similar to existing techniques for bounding the Fiedler value $\lambda_{2}(G)$ of graphs in these classes [ST96, ST07, Kel04, Kel06b, Kel06a]. At a high-level, the method is always to map a set of $d$-dimensional balls representing the graph to a $d$-dimensional (hyper)sphere (that is, the (hyper)surface of a $(d+1)$-dimensional ball) so that the measure of the surface of the (hyper)sphere can be related to the average ply of the balls. In most cases, it is easier to bound the maximum ply of the balls, but genus-$g$ graphs require a more robust approach.
2.3.1 New techniques for bounded-genus graphs
While most of these bounds can be obtained primarily by adapting known techniques, we require some very significant novel constructions to prove Proposition 2.7.
First, we show that Kelner’s method for triangulated genus-$g$ graphs that obtains a bound of $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$ can also be extended to bound $\gamma^{(1)}(G)$. Kelner’s method is quite involved, so there are many aspects that have to be re-checked in the context of our modified problem, but this step uses similar methods overall.
However, this is not enough to obtain our bound of $\gamma^{(n)}(G)\lesssim\frac{g\cdot\log\Delta}{n}$, which applies not just to triangulated genus-$g$ graphs but also to general genus-$g$ graphs. To obtain this bound, we provide a pair of results:
Lemma 2.8.
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices, maximum degree $4$, and genus $g$, so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
Lemma 2.9.
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
Both of these results use a new structure we call uniform shallow minors, which are essentially shallow minors in which all collections of vertices are of exactly equal size, and all vertices are in some collection. One could also think of these as a very special kind of bounded-diameter decomposition. However, the way we use this structure is in some sense opposite to how both minors and bounded-diameter decompositions are typically used: In each result, we construct a graph $H$ that contains $G$ as a uniform shallow minor.
2.3.2 A new separator theorem for ball-intersection graphs of bounded ply
For the case of $d$-dimensional ball-intersection graphs of ply $k$, and $d$-dimensional $k$-nearest-neighbour graphs, the sizes of the separators guaranteed to exist by Corollary 1.6 are actually improvements over the previous bests by a factor of $\frac{\sqrt{\min\{d,\log\Delta\}}}{d}$. The previous best bound for both classes was a proof that $\operatorname{\mathcal{O}}(d)$-balanced vertex-separators exist of size $\operatorname{\mathcal{O}}(k^{1/d}n^{1-1/d})$ [MTTV97], which implies the existence of a $\frac{2}{3}$-balanced vertex separator of size $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$. While our separators can’t be constructed via reweighted spectral partitioning, they still can be constructed in polynomial time if the explicit geometric representation of the graph is given. We will obtain the following results:
Corollary 2.10.
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
Corollary 2.11.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
2.4 Bounds via Extremal Spread
Our second family of methods for obtaining upper bounds on $\gamma^{(n)}$ and $\gamma^{(1)}$ makes use of a widely-used quantity for approximation algorithms: Extremal spread of (vertex-weighted) shortest-path metrics. Intuitively, extremal spread is a (normalized) weighting of the vertices in a graph so that the pairwise distances are maximized.
@@ 2.4 Bounds via Extremal Spread
$L^{1}$-extremal spread is known to be closely related to vertex expansion. However, in this work, we are exclusively interested in the case of $L^{2}$-extremal spread, which has two very important known bounds:
Proposition 2.13 ([BLR10, Theorem 3.1]).
For a genus-$g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
@@
Proposition 2.14 ([BLR10, DP25]).
For a $K_{h}$-minor-free graph $G$ on $n$ vertices, if $n\gtrsim h\log\log h$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h\log\log h}.$
The first of these theorems is optimal up to constant factors, while the second is optimal up to the $\operatorname{\mathcal{O}}(\log\log h)$ factors. These $\operatorname{\mathcal{O}}(\log\log h)$ factors come from the current state of progress towards Hadwiger’s conjecture, a major open problem in graph theory [Die25].
The key result we will obtain is as follows:
Proposition 2.15.
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(1)}(G)\lesssim\left[\alpha(G)\right]^{2}\cdot\frac{n^{3}}{\left[\overline{s}_{2}(G)\right]^{2}}.$
3 The Algorithm
We will now describe the three steps of the reweighted spectral partitioning algorithm in detail. This is only a reference for the algorithm and its time complexity/variations – the corresponding theorems are stated in the next section.
We start by presenting the simplest possible form of the algorithm, which is Monte Carlo randomized.
Dimension Reduction
We obtain a solution to $\gamma^{(1)}$ from the solution to $\gamma^{(n)}$.
Algorithm 1 ProjectionReduce($f:V\to\mathbb{R}^{n}$, $G=(V,E)$)
for $i=1$ to $\Theta(\log n)$ do
  Sample $r_{i}\sim\mathcal{N}(0,1)^{n}$
  $f^{\prime}_{i}(v)\leftarrow\langle f(v),r_{i}\rangle$ for all $v\in V$
end for
return $f^{\prime}_{i}$ maximizing $\sum_{u,v\in V}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
Algorithm 2 DistanceReduce($y:V\to\mathbb{R}$, $G=(V,E)$, RandomPartition)
Compute $d(u,v)$ for all $u,v\in V$ as shortest-path distance in $G$ with vertex weights $y$
$S_{0}\leftarrow\left\{\arg\max_{v\in V}\sum_{u\in V}d(u,v)^{2}\right\}$
for $i=1$ to $\Theta(\log n)$ do
  $\mathcal{P}\leftarrow$ RandomPartition(G,y)(G,y) $\triangleright$ $\mathcal{P}$ is a partition of the vertices
  Let $S_{i}$ be the union of parts $P\in\mathcal{P}$ each sampled with probability $\frac{1}{2}$
end for
for each $S_{i}$ do
  $f^{\prime}_{i}(u)\leftarrow\min_{v\in S_{i}}d(u,v)$ for all $u\in V$
end for
return $f^{\prime}_{i}$ maximizing $\sum_{uv\in E}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
Algorithm 3 BestReduce($f:V\to\mathbb{R}^{n}$, $y:V\to\mathbb{R}$, $G$, RandomPartition $\bot$)
$f^{\prime}_{1}\leftarrow$ ProjectionReduce($f$, $G$)
if RandomPartition $\neq\bot$ then $\triangleright$ If optional partition oracle is provided, use it
  $f^{\prime}_{2}\leftarrow$ DistanceReduce($y$, $G$, RandomPartition)
  return $f^{\prime}_{i}$ maximizing $\sum_{uv\in E}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
else
  return $f^{\prime}_{1}$
end if
Sweep Algorithm
Extract a set of vertices from the one-dimensional embedding.
Algorithm 4 Sweep($f^{\prime}:V\to\mathbb{R}$, $y:V\to\mathbb{R}$, $G=(V,E)$)
Sort vertices by $f^{\prime}$ value: $v_{1},\dots,v_{n}$
Initialize current set of crossing edges $\mathcal{C}\leftarrow\emptyset$
for $k=1$ to n1n-1 do
  $A_{k}\leftarrow\{v_{1},\dots,v_{k}\}$, $B_{k}\leftarrow\{v_{k+1},\dots,v_{n}\}$
  Update $\mathcal{C}$ by adding edges in $v_{k}\times B_{k}\cap E$ and removing edges in $v_{k}\times A_{k}\cap E$.
  $S_{k}\leftarrow\{u:uv\in\mathcal{C},y(u)\geq y(v)\}\cup\{v:uv\in\mathcal{C},y(v)>y(u)\}$
  $\psi_{k}\leftarrow|S_{k}|/\min\{|A_{k}|,|B_{k}|\}$
end for
return the partition $(A_{k},B_{k},S_{k})$ minimizing $\psi_{k}$
The Complete Algorithms
We can use one pass of all three steps to obtain a cut with small vertex expansion, or use $\operatorname{\mathcal{O}}(n)$ passes to obtain a balanced vertex separator.
Algorithm 5 ReweightedSpectralPartitioning($G=(V,E)$, RandomPartition $\bot$)
Step 1: Solve the SDP $\gamma^{(n)}(G)$, obtaining $f:V\to\mathbb{R}^{n}$ and $y:V\to\mathbb{R}$
Step 2: $f^{\prime}\leftarrow$ BestReduce($f$, $y$, $G$, RandomPartition)
Step 3: $(A,B,S)\leftarrow$ Sweep($f^{\prime}$, $y$, $G$)
return $(A,B,S)$
Algorithm 6 BalancedRSP($G=(V,E)$, RandomPartition $\bot$)
Initialize $S_{\mathrm{union}}\leftarrow\emptyset$
$V_{\mathrm{residual}}\leftarrow V$
while $|V_{\mathrm{residual}}|>\frac{2}{3}|V|$ do
  Let $G_{\mathrm{residual}}$ be the induced subgraph of $G$ on $V_{\mathrm{residual}}$
  $(A,B,S)\leftarrow$ ReweightedSpectralPartitioning($G_{\mathrm{residual}}$, RandomPartition)
  $S_{\mathrm{union}}\leftarrow S_{\mathrm{union}}\cup S$
  $V_{\mathrm{residual}}\leftarrow\arg\max_{X\in\{A,B\}}|X|$
end while
return $S_{\mathrm{union}}$
Note that Algorithm 6 is exactly the algorithm one can apply to get the guarantees in Corollary 1.7 and Corollary 1.8.
3.1 Practical considerations
As previously mentioned, we have implemented this algorithm (specifically, Algorithm 6), although without support for partition oracles. To solve the SDP, we used the Clarabel solver [GC24]. Unfortunately, practical implementations for general SDPs are still quite slow and memory-inefficient (they often involve computing the Hessian), and this was not able to compute $\gamma^{(n)}$ for even planar graphs beyond around 150 vertices. Boyd, Diaconis, and Xiao described and tested a much more practical projected subgradient method for this problem [BDX04], but it works in the formulation of $\lambda_{2}^{*}$ rather than $\gamma^{(n)}$, and so does not natively provide the vectors $f(v)\in\mathbb{R}^{n}$ that we use for dimension-reduction and rounding. Moreover, it has no clear stopping criterion to guarantee any level of optimality, and it is not guaranteed to run in polynomial time.
4 A Refined Cheeger Inequality
In this section, we will prove Theorem 1.4 via the algorithm outlined in Section 3. Fix a graph $G=(V,E)$ with maximum degree $\Delta$. Several of the inequalities stated in Theorem 1.4 are already known:
Theorem 4.1 ([OTZ22, Theorem 2.10]).
For a graph $G$,
$\psi(G)^{2}\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
The result that $\psi(G)^{2}\lesssim\gamma^{(1)}(G)$ was re-proven by Kwok, Lau, and Tung [KLT22], who gave a slightly simpler algorithm that takes the best of a set of candidate cuts via a sweep. Our Algorithm 4 further simplifies their algorithm and considers a superset of the cuts that they consider.
Theorem 4.2 ([KLT22, Proposition 3.14]).
For a graph $G$ with maximum degree $\Delta$,
The proof that $\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot\log\Delta$ uses a random projection and succeeds in expectation, so Algorithm 1 replicates this guarantee with high probability.
Thus, the only remaining inequality in Theorem 1.4 left for us to prove is:
$\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot[\alpha(G)]^{2}.$
In order to prove the result, we must also do so in a way that shows Algorithm 2 produces a feasible solution for $\gamma^{(1)}(G)$ obtaining this bound with high probability.
We will do this in several steps. First, we define the following quantity: